OpenStudy (anonymous):

The volume of a cylinder is one hundred twenty cubic meters. What is the volume of a cone with the same radius and the same height as the cylinder? thirty cubic meters forty cubic meters three hundred sixty cubic meters four hundred eighty cubic meters

OpenStudy (anonymous):

Please help.

OpenStudy (chrisisdope1):

B because you can fit in 3 cones in one cylinder.

OpenStudy (tkhunny):

Right Circular Cylinder Volume = \(\pi r^{2}h\) Right Circular Cone Volume = \(\dfrac{1}{3}\pi r^{2}h\)

OpenStudy (chrisisdope1):

And the volume of cone is 1/3 rd volume of cylindr with same dimensions..

OpenStudy (help!!!!):

use \[v=pir ^{2}h\]

OpenStudy (anonymous):

Thank you so much